# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

### v3

@@ 2. Tall local models
Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$. Let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ and $\Phi_{K}\colon{\mathbb{C}}^{k+1}\to\mathfrak{k}^{*}$ be the homogeneous moment maps for the $H$-action on ${\mathbb{C}}^{h+1}$ and the $K$-action on ${\mathbb{C}}^{k+1}$, respectively. Since $\xi_{i}=0$ for all $i>k$, we have $\Phi_{H}^{-1}(0)=\Phi_{K}^{-1}(0)\times\{0\}$, and so $\Phi_{Y}^{-1}(0)=T\times_{H}\{0\}\times(\Phi_{K}^{-1}(0)\times\{0\})$, where $\Phi_{Y}:Y\to{\mathfrak{t}}^{*}$ is the moment map for the $T$-action on $Y$. Since $\rho(H)=K\times(S^{1})^{h-k}$, the claim follows immediately. ∎
On the reduced space Y//T{Y/\!/T} there is a simple relation between the magnitudes of the defining polynomial and of $z$.
@@
Let $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}$. Let $N$ be the degree of the defining polynomial $P(z)=\prod_{j}z_{j}^{\xi_{j}}$. Then for all $t,\alpha,z]\in{Y/\!/T}$,
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model. By a slight abuse of notation, we call a $T$-invariant function $g\colon Y\to{\mathbb{R}}$ a $\boldsymbol{T}$-invariant (homogeneous) polynomial if its restriction to ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a (homogeneous) polynomial.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The algebra of $T$-invariant polynomials on $Y$ is generated by the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the map $t,\alpha,z]\mapsto|z|^{2}$, and the components of $\Phi_{Y}$.
@@
By definition, there is a one-to-one correspondence between $T$-invariant polynomials on $Y$ and $H$-invariant polynomials on ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$. Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$ and let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ be the homogeneous moment map. Since $\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z)$ for all $t,\alpha,z]\in Y$, it suffices to prove that the algebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ is generated by the real and imaginary parts of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$, the map $z\mapsto|z|^{2}$, and the components of $\Phi_{H}$. Since the local model $Y$ is tall, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$. Hence, since $\Phi_{H}(z)=\frac{1}{2}\rho^{*}(|z_{0}|^{2},\ldots,|z_{h}|^{2})$, the subalgebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ generated by the map $z\mapsto|z|^{2}$ and the components of $\Phi_{H}$ is also generated by the functions $z\mapsto|z_{i}|^{2}$ for $i=0,\ldots,h$.
@@
The proof of Lemma 2.11 provides an equivalent generating set for the algebra of $T$-invariant polynomials on a tall local model $Y$, namely the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the maps $|z_{i}|^{2}\colon Y\to{\mathbb{R}}$ that send $t,\alpha,z]$ to $|z_{i}|^{2}$ for $i=0,\ldots,h$, and the components of the map $t,\alpha,z]\mapsto\alpha$.
@@ 3. The key proposition
Near each regular point of $g$ modulo $\Phi$ there is a natural structure of smooth manifold on the reduced space so that $\overline{g}$ is smooth with no critical points. More precisely, the following holds.
@@
Moreover, $p$ is tall, has purely elliptic type, and the defining polynomial of the local model for $p$ has degree $1$.
@@
The result is local. Hence, by the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. First we show that $Y$ is tall. If it is short, then $\Phi^{-1}(\Phi(p))=T\cdot p$. Since $\Phi$ is invariant under the Hamiltonian flow of $g$, this implies that $(T\times{\mathbb{R}})\cdot p=T\cdot p$. Hence $p$ is a critical point of $g$ modulo $\Phi$, which contradicts our assumption.
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is given by ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. By Remark 2.12, if the degree $N$ of the defining polynomial $P$ is greater than 1, then the only $T$-invariant polynomials of degree 1 on $Y$ are the components of the map $t,\alpha,z]\mapsto\alpha$, while if $N=1$, the real and imaginary parts of $P$ are also $T$-invariant polynomials of degree 1. Since the derivative of the map $t,\alpha,z]\mapsto\alpha$ at $p$ vanishes when restricted to ${\operatorname{ker}}D_{p}\Phi$, this implies that $N=1$. If we assume, without loss of generality, that $P(z)=z_{0}$, it also implies that $D_{p}g$ restricts to a non-zero $H$-invariant linear function on $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$. By Lemma 2.8, $\overline{P}$ is a diffeomorphism and $\overline{g}\circ\overline{P}^{-1}\colon{\mathbb{C}}\simeq{\mathbb{R}}^{2}\to{\mathbb{R}}$ is equal to the restriction of $g$ to ${\mathbb{C}}\simeq\{1\}\times_{H}\{0\}\times({\mathbb{C}}\times\{0\})\subseteq Y$. Hence, since the restriction of $D_{p}g$ to $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$ is non-zero, there exists an open neighborhood $U_{p}$ of $[p]$ in Y//T{Y/\!/T} such that the restriction of $\overline{g}\circ\overline{P}^{-1}$ to $\overline{P}(U_{p})$ is smooth and has no critical points. Setting $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ to be the restriction of $\overline{P}$ to $U_{p}$, the first result follows.
@@
Clearly, every critical point of $g$ modulo $\Phi$ is a singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$. The converse holds over the interior of $\Phi(M)$ if $\Phi$ is proper. To see this, let $p$ be a regular point of $g$ modulo $\Phi$. By Lemma 3.2, $p$ is tall and the degree of the defining polynomial of the local model for $p$ is $1$. Moreover, if $\Phi(p)$ is in the interior of $\Phi(M)$, then the local model for $p$ has a surjective moment map because $\Phi$ is open as a map to $\Phi(M)$ by Theorem 5.5. Together, these facts imply that the $T$-stabilizer of $p$ is trivial. Therefore, $p$ is a regular point of $\Phi$, and so also of $f$.
@@
Assume for simplicity that $\beta=0$. By hypothesis and Lemma 3.2, for each $p\in\Phi^{-1}(0)$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space ${M/\!/T}:=\Phi^{-1}(0)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above. Since M//T{M/\!/T} is paracompact, the open cover $\{U_{p}\}_{p\in\Phi^{-1}(0)}$ has a locally finite refinement. Hence, there exists a proper subset $\mathcal{P}\varsubsetneq\Phi^{-1}(0)$ and a locally finite open cover $\{U_{p}\}_{p\in\mathcal{P}}$ such that there exists a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above for each $p\in\mathcal{P}$; however, we may no longer assume that $U_{p}$ contains $p$. Since a locally finite set of points is closed, we may also assume that if $p]\in U_{p}$ then $p]\notin U_{p^{\prime}}$ for all $p^{\prime}\neq p$ in $\mathcal{P}$. By construction, for all $p,p^{\prime}\in\mathcal{P}$, the transition function $\Psi_{p}\circ\Psi_{p^{\prime}}^{-1}$ is a diffeomorphism from $\Psi_{p^{\prime}}(U_{p}\cap U_{p^{\prime}})$ to $\Psi_{p}(U_{p}\cap U_{p^{\prime}})$. Hence, the coordinate charts $\Psi_{p}$ endow the reduced space M//T{M/\!/T} with the structure of a $2$-dimensional smooth manifold. Moreover, with respect to this smooth structure, $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function. Additionally, $[p]$ is a critical point of $\overline{g}$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$; in this case, the index of $[p]$ is the index of $\Psi_{p}([p])$. Finally, since every point in $\Phi^{-1}(0)$ is tall, M//T{M/\!/T} is oriented by [KT1, Proposition 6.1]. ∎

### v4

@@ 2. Tall local models
Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$. Let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ and $\Phi_{K}\colon{\mathbb{C}}^{k+1}\to\mathfrak{k}^{*}$ be the homogeneous moment maps for the $H$-action on ${\mathbb{C}}^{h+1}$ and the $K$-action on ${\mathbb{C}}^{k+1}$, respectively. Since $\xi_{i}=0$ for all $i>k$, we have $\Phi_{H}^{-1}(0)=\Phi_{K}^{-1}(0)\times\{0\}$, and so $\Phi_{Y}^{-1}(0)=T\times_{H}(\{0\}\times(\Phi_{K}^{-1}(0)\times\{0\}))$, where $\Phi_{Y}:Y\to{\mathfrak{t}}^{*}$ is the moment map for the $T$-action on $Y$. Since $\rho(H)=K\times(S^{1})^{h-k}$, the claim follows immediately. ∎
On the reduced space Y//T{Y/\!/T} there is a simple relation between the magnitudes of the defining monomial and of $z$.
@@
Let $Y:=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a tall local model with moment map $\Phi_{Y}$. Let $N$ be the degree of the defining monomial $P(z)=\prod_{j}z_{j}^{\xi_{j}}$. Then for all $t,\alpha,z]\in{Y/\!/T}$,
@@
Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a local model. By a slight abuse of notation, we call a $T$-invariant function $g\colon Y\to{\mathbb{R}}$ a $\boldsymbol{T}$-invariant (homogeneous) polynomial if its restriction to ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a (homogeneous) polynomial.
@@
Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The algebra of $T$-invariant polynomials on $Y$ is generated by the real and imaginary parts of the defining monomial $P\colon Y\to{\mathbb{C}}$, the map $t,\alpha,z]\mapsto|z|^{2}$, and the components of $\Phi_{Y}$.
@@
By definition, there is a one-to-one correspondence between $T$-invariant polynomials on $Y$ and $H$-invariant polynomials on ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$. Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$ and let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ be the homogeneous moment map. Since $\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z)$ for all $t,\alpha,z]\in Y$, it suffices to prove that the algebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ is generated by the real and imaginary parts of the defining monomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$, the map $z\mapsto|z|^{2}$, and the components of $\Phi_{H}$. Since the local model $Y$ is tall, Remark 2.5 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$. Hence, since $\Phi_{H}(z)=\frac{1}{2}\rho^{*}(|z_{0}|^{2},\ldots,|z_{h}|^{2})$, the subalgebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ generated by the map $z\mapsto|z|^{2}$ and the components of $\Phi_{H}$ is also generated by the functions $z\mapsto|z_{i}|^{2}$ for $i=0,\ldots,h$.
@@
The proof of Lemma 2.14 provides an equivalent generating set for the algebra of $T$-invariant polynomials on a tall local model $Y$, namely the real and imaginary parts of the defining monomial $P\colon Y\to{\mathbb{C}}$, the maps $|z_{i}|^{2}\colon Y\to{\mathbb{R}}$ that send $t,\alpha,z]$ to $|z_{i}|^{2}$ for $i=0,\ldots,h$, and the components of the map $t,\alpha,z]\mapsto\alpha$.
@@ 3. The key proposition
Clearly, every critical point of $g$ modulo $\Phi$ is a singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$. If $\Phi$ is proper then the converse holds over the interior of $\Phi(M)$; see Remark 3.5 below.
Remark 3.2.
By linear algebra, $({\operatorname{ker}}D_{p}\Phi)^{\circ}=\{D_{p}\Phi^{\xi}\mid\xi\in{\mathfrak{t}}\}$. Hence, $p$ is a so-called “relative equilibrium of the Hamiltonian vector field $X_{g}$ with respect to the symmetry given by the $T$-action,” i.e., there exists $\xi\in{\mathfrak{t}}$ such that $D_{p}g=D_{p}\Phi^{\xi}$, exactly if $D_{p}g\in({\operatorname{ker}}D_{p}\Phi)^{\circ}$, which occurs exactly if $p$ is a critical point of $g$ modulo $\Phi$.
To understand the next lemma, recall that by Lemma 2.10 there is a natural structure of smooth manifold on the reduced spacenear each non-exceptional orbit, with respect to which the induced function $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is smooth.
Lemma 3.3.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. If $p\in\Phi^{-1}(0)$ is not exceptional, then $[p]$ is a critical point of $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ exactly if $p$ is critical point of $g$ modulo $\Phi$.
Proof.
By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Since $p$ is not exceptional, $p$ is tall. Let $P(z)=\prod_{i=0}^{h}z_{i}^{\xi_{i}}$ be the defining monomial of $Y$. Since $\sum_{i}\xi_{i}=1$ by Lemma 2.9, we may further assume that ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ and $H\simeq(S^{1})^{h}$, where $(S^{1})^{h}$ acts on ${\mathbb{C}}^{h}$ in the standard fashion and trivially on $\mathbb{C}$}.
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. Moreover, since $g$ is $T$-invariant and $H$ acts on ${\mathbb{C}}^{h}$ via an isomorphism with $(S^{1})^{h}$, the functional $D_{p}g$ is $H$-invariant and vanishes on $\{0\}\times\{0\}\times(\{0\}\times{\mathbb{C}}^{h})$. Hence, $p$ is a critical point of $g$ modulo $\Phi$ exactly if $D_{p}g$ vanishes on ${\mathfrak{h}}^{\perp}\times\{0\}\times({\mathbb{C}}\times\{0\})$. Since $\Phi^{-1}(0)=T\times_{H}(\{0\}\times({\mathbb{C}}\times\{0\})$, this holds exactly if $[p]$ is a critical point of $\overline{g}$. ∎
There is also a natural structure of a smooth manifold on the reduced space near each regular point of $g$ modulo $\Phi$, but in this case the induced function $\overline{g}$ is locally regular. More precisely, the following holds.
@@
Moreover, $p$ is tall, has purely elliptic type, and the defining monomial of the local model for $p$ has degree $1$.
@@
By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. First we show that $Y$ is tall. If it is short, then $\Phi^{-1}(\Phi(p))=T\cdot p$. Since $\Phi$ is invariant under the Hamiltonian flow of $g$, this implies that $(T\times{\mathbb{R}})\cdot p=T\cdot p$. Hence $p$ is a critical point of $g$ modulo $\Phi$, which contradicts our assumption.
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is given by ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. By Remark 2.15, if the degree $N$ of the defining monomial $P$ is greater than 1, then the only $T$-invariant polynomials of degree 1 on $Y$ are the components of the map $t,\alpha,z]\mapsto\alpha$, while if $N=1$, the real and imaginary parts of $P$ are also $T$-invariant polynomials of degree 1. Since the derivative of the map $t,\alpha,z]\mapsto\alpha$ at $p$ vanishes when restricted to ${\operatorname{ker}}D_{p}\Phi$, this implies that $N=1$. If we assume, without loss of generality, that $P(z)=z_{0}$, it also implies that $D_{p}g$ restricts to a non-zero $H$-invariant linear function on $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$. By Lemmas 2.10, $\overline{P}$ is a diffeomorphism; its inverse is given by $z\mapsto[1,0,(z,0)]$. Therefore $\overline{g}\circ\overline{P}^{-1}\colon{\mathbb{C}}\simeq{\mathbb{R}}^{2}\to{\mathbb{R}}$ is equal to the restriction of $g$ to ${\mathbb{C}}\simeq\{1\}\times_{H}(\{0\}\times({\mathbb{C}}\times\{0\}))\subseteq Y$. Hence, since the restriction of $D_{p}g$ to $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$ is non-zero, there exists an open neighborhood $U_{p}$ of $[p]$ in Y//T{Y/\!/T} such that the restriction of $\overline{g}\circ\overline{P}^{-1}$ to $\overline{P}(U_{p})$ is smooth and has no critical points. Setting $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ to be the restriction of $\overline{P}$ to $U_{p}$, the first result follows.
@@
If $\Phi$ is proper, then every singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ over the interior of $\Phi(M)$ is a critical point of $g$ modulo $\Phi$. To see this, let $p$ be a regular point of $g$ modulo $\Phi$. By Lemma 3.4, $p$ is tall and the degree of the defining monomial of the local model for $p$ is $1$. Moreover, if $\Phi(p)$ is in the interior of $\Phi(M)$, then the local model for $p$ has a surjective moment map because $\Phi$ is open as a map to $\Phi(M)$ by Theorem 5.5. Together, these facts imply that the $T$-stabilizer of $p$ is trivial. Therefore, $p$ is a regular point of $\Phi$, and so also of $f$.
@@
Assume for simplicity that $\beta=0$. By hypothesis and Lemma 3.4, for each $p\in\Phi^{-1}(0)$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space ${M/\!/T}:=\Phi^{-1}(0)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above; indeed, if $p$ is a regular point of $g$ modulo $\Phi$, then the stronger properties (a) and (b) of Lemma 3.3 also hold. In particular, $\Psi_{p}$ is a diffeomorphism on the subset of orbits of regular points of $g$ modulo $\Phi$. By shrinking $U_{p}$ if necessary, we may further assume that $\overline{g}\circ\Psi_{p}^{-1}$ is a regular function on $\Psi_{p}(U_{p}\smallsetminus\{[p]\})$. Moreover, by the local normal form and Lemma 2.10, the set of exceptional orbits in M//T{M/\!/T} is discrete. Therefore, by further shrinking $U_{p}$ if necessary, we may assume that $U_{p}\smallsetminus\{[p]\}$ contains no exceptional orbits. By Lemma 3.3, this implies that $U_{p}\smallsetminus\{[p]\}$ contains no critical points of $g$ modulo $\Phi$.
Hence we may choose a subset $\mathcal{P}\varsubsetneq\Phi^{-1}(0)$ such that $\{U_{p}\}_{p\in\mathcal{P}}$ is a cover of M//T{M/\!/T} and each orbit of critical points of $g$ modulo $\Phi$ is contained in exactly one open set. Therefore for all $p,p^{\prime}\in\mathcal{P}$, the transition function $\Psi_{p}\circ\Psi_{p^{\prime}}^{-1}$ is a diffeomorphism from $\Psi_{p^{\prime}}(U_{p}\cap U_{p^{\prime}})$ to $\Psi_{p}(U_{p}\cap U_{p^{\prime}})$. Hence, the coordinate charts $\Psi_{p}$ endow the reduced space M//T{M/\!/T} with the structure of a $2$-dimensional smooth manifold. Moreover, with respect to this smooth structure, $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function. Additionally, $[p]$ is a critical point of $\overline{g}$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$; in this case, the index of $[p]$ is the index of $\Psi_{p}([p])$. Finally, since every point in $\Phi^{-1}(0)$ is tall, M//T{M/\!/T} is oriented by [KT1, Proposition 6.1]. ∎